The thrust of the XLR99 rocket engine was essentially constant at 57,000 lb during the powered phase of the X-15 flight. For the maximum altitude flight on August 22,1963, the aerodynamic drag at the burnout altitude of 176,000 ft was only 252 lb, due to very low air density at that altitude. The empty weight of the X-15 was 14,590 lb. Using the equation of rate of climb l, which contains the assumption of steady unaccelerated flight, calculate the rate if climb of the X-15


1
Expert's answer
2021-06-15T13:56:01-0400

"ROC =(\frac{T-D}{W})V=(\frac{57000-252}{14590})*5600=21781.27 ft/s"